The all-equity firm Metallica Corporation wants to diversify its operations. Some recent financial information for the company is shown here:

  Stock price $ 74
  Number of shares 30,000
  Total assets $ 9,800,000
  Total liabilities $ 4,700,000
  Net income $ 420,000

MHMM is considering an investment that has the same PE ratio as the firm. The cost of the investment is $640,000, and it will be financed with a new equity issue. The return on the investment will equal MHMM’s current ROE

What is the new price per share if the investment is made?

What is the new market to book ratio?

What is the new earnings per share?

What is the NPV of this investment?
Answer #1

Current Equity of Company =Total Assets - Total Liabilities

=9,800,000-4,700,000

=$5100000

Return on Equity = Net Income/Total Equity

=420000/5100000

=8.24%

New total Equity ie total equity after investment =$5100000+640000=$5740000

Net Income after investment =$5740000*8.24%=$427606

The company’s current earnings per share ie before investment =Net income/no of shares =420000/30000=14 per share

Number of new shares = Investment/ Current Market price 640000/$74=8649 shares

The earnings per share after the stock offer (new earnings per share) = net income /No of shares = 427606/(30000+8649)=11.06 per share

The current P/E ratio is:price/EPS =74/14=5.29

What is the new price per share if the investment is made?

Hence price of share will be P/E ratio * EPS =5.29*11.06=58.48 per Share

What is the new market to book ratio?

The current book value per share =Equity/No of Shares =5100000/30000 =170 per Share

Book value share after investment =(5100000+640000)/38649=$148.52

Market-to-book ratio= $58.48/$148.52=58.48/148.52=39.38%

The NPV of the project = the cost of the project + the new market value of the firm - the current market value of the firm

=640000+38649*58.48-74*30000

NPV=$680193